OpenStudy (fanduekisses):

let f(x)=g(x)h(x), g(10)=-4, h(10)=560, g'(10)=0, and h'(10)=35. Find f'(10)

OpenStudy (fanduekisses):

Do I have to use the slope of the tangent line?

OpenStudy (tkhunny):

Have you considered the Product Rule? f'(x) = g(x)h'(x) + g'(x)h(x) Should be familiar.

OpenStudy (fanduekisses):

Ohhh yeah

OpenStudy (fanduekisses):

Let me try it :)

OpenStudy (fanduekisses):

(-4)(35)+(560)(0) =140

OpenStudy (tkhunny):

f'(x) = g(x)h'(x) + g'(x)h(x) g(10)=-4, h(10)=560, g'(10)=0, and h'(10)=35 f'(10) = (-4)(35) + (0)h(x) = -140 You were SO CLOSE!!

OpenStudy (fanduekisses):

oops forgot about the (-) sign hehe thanks ^_^

OpenStudy (tkhunny):

Notice how I didn't bother to substitute for h(10). No matter what it is, the product woudl be zero, so I stopped.